OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (texaschic101):

First we have to find the slope of the equation given. The reason we do this is because a parallel line will have the same slope. Remember that in y = mx + b form, the number in the m position represents the slope. So we will put our equation in y = mx + b form. 3x - y = -3 --- subtract 3x from both sides -y = -3x - 3 --- multiply by -1 to make y positive y = 3x + 3 The number in the m position is 3, therefore, that is our slope. Now we will use y = mx + b form again. y = mx + b we already know the slope(m) = 3 (4,-2).....x = 4 and y = -2 now we just need to find b, the y intercept so lets sub -2 = 3(4) + b -2 = 12 + b -2 - 12 = b -14 = b so our parallel equation is : y = 3x - 14